1 upon x + 1 upon 3x - 1 upon 5x
Answer by stanbon(75887)

1 upon x + 1 upon 3x - 1 upon 5x
---
(1/x) + 1/(3x) - (1/5xO
---
least common denominator: 15x
---
= 15/lcd + 5/lcd - 3/lcd
-----
= (15 + 5 -3)/lcd
-----
= (17)/15x
==============
Cheers,
Stan H.